OpenStudy (anonymous):

A rectangle has a length of the cube root of 81 inches and a width of 3 to the 2 over 3 power inches. Find the area of the rectangle.

OpenStudy (danjs):

It will be the product of the two...

OpenStudy (danjs):

\[l = \sqrt[3]{81} = 81^{1/3}\] \[w = 3^{2/3} = [3^2]^{1/3} = 9^{1/3}\]

OpenStudy (danjs):

\[A = l*w = 81^{1/3} * 9^{1/3} = [{81*9}]^{1/3} = 729^{1/3} = 9\]

OpenStudy (danjs):

just a bunch of power rules